[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 20, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 21 02:46:17 UTC 2016


HSU Shift Log for Wed Jul 20 2016 22:46:12 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: The Pacific ATCF was very slow earlier, but improved some as  
the night went on. Also the microwave data was not coming in for Estelle,  
but the problem seems to have resolved itself by 23Z

N-AWIPS/AWIPS2 issues: Ridges has had a lot of activity even when I was  
doing very little with the machine-- the load bar on the lower left has  
been very high. I logged in and logged off and the load remains high. It  
has decreased since earlier, however.

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160721/dc0f41f8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list